# ORBIT: On-policy Exploration-Exploitation for Controllable Multi-Budget Reasoning

### v1

@@ Abstract
Recent Large Reasoning Models (LRMs) achieve strong performance by leveraging long-form Chain-of-Thought (CoT) reasoning, but uniformly applying overlong reasoning at inference time incurs substantial and often unnecessary computational cost. To address this, prior work explores various strategies to infer an appropriate reasoning budget from the input. However, such approaches are unreliable in the worst case, as estimating the minimal required reasoning effort is fundamentally difficult, and they implicitly fix the trade-off between reasoning cost and accuracy during training, limiting flexibility under varying deployment scenarios. Motivated by these limitations, we propose ORBIT, a controllable multi-budget reasoning framework with well-separated reasoning modes triggered by input. ORBIT employs multi-stage reinforcement learning to discover Pareto-optimal reasoning behaviors at each effort, followed by on-policy distillation to fuse these behaviors into a single unified model. Experiments show that ORBIT achieves (1) controllable reasoning behavior over multiple modes, (2) competitive reasoning density within each mode, and (3) integration of these frontier policies into a single unified student model while preserving clear mode separation and high per-mode performance.
@@ 1 Introduction
Figure 1: ORBIT demonstrates controllable reasoning behavior comparable to advanced systems, covering multiple reasoning modes: low/mid/high for o3-mini and gpt-oss-120B, and an extended set of low/mid/high/extra-high for ORBIT.
Recent advances in Reinforcement Learning with Verifiable Rewards (RLVR) (Guo et al., 2025; Team et al., 2025; Xiaomi et al., 2025) have significantly pushed the frontier of complex task solving and long-form Chain-of-Thought (CoT) (Wei et al., 2022) reasoning in large language models. By encouraging explicit reasoning trajectories, these methods enable models to explore richer solution paths and achieve strong performance on challenging benchmarks (Guo et al., 2025). However, the continued push toward greater reasoning depth introduces a practical trade-off between reasoning effort and operational efficiency: increasingly long reasoning traces incur higher latency and computational cost, and are often unnecessary for simpler queries (Yang et al., 2025b).
Notably, this observation concerns not the use of long-chain reasoning during training, but its uncalibrated application at inference time, as longer reasoning chains often facilitate exploration of richer solution paths and contribute to higher peak accuracy (Setlur et al., 2025). The challenge primarily emerges at deployment time, where models trained to reason deeply may apply the same heavy reasoning patterns indiscriminately across inputs.
To optimize inference-time efficiency, prior work has explored various strategies for adjusting reasoning effort according to task demands. (Luo et al., 2025a; Lou et al., 2025; Luo et al., 2025b; Zhang et al., 2025). Across these approaches, the underlying goal is effectively the same: to estimate an optimal allocation of reasoning effort.
However, this estimation problem is fundamentally unreliable. Determining the minimal amount of reasoning required to solve an instance is closely related to estimating its Kolmogorov complexity (Li and Vitnyi, 2008), which is known to be uncomputable and connected to the halting problem theoretically (Graves, 2017). Empirically, adaptive mechanisms can also misallocate reasoning effort—truncating reasoning on instances that would benefit from deeper deliberation (Aggarwal et al., 2025; Jin et al., 2025), while preserving unnecessarily long reasoning for simpler instances (Chen et al., 2025a).
@@
Motivated by these limitations, we rethink a controllable reasoning model that can flexibly adjust its reasoning effort according to user preferences at inference time, meanwhile achieving competitive reasoning densities within each effort.
We introduce ORBIT, a fully on-policy training framework for controllable multi-budget reasoning. By exploring under different training context windows, ORBIT identifies a Pareto frontier of reasoning behaviors and effectively integrates these into a single model whose discrete reasoning modes can be triggered by prompts. As illustrated in Figure 1, ORBIT achieves clear and well-separated reasoning modes similar to state-of-the-art controllable systems, while maintaining strong per-mode accuracy and token-efficient reasoning.
@@
    We propose ORBIT, a fully on-policy training framework for controllable multi-budget reasoning, enabling a single model to generate responses across multiple reasoning modes triggered by mode-specific queries.
@@
    We introduce an expansion-compression exploration strategy that efficiently discovers mode-specific reasoning frontiers under varying context constraints, yielding high-accuracy policies at each reasoning effort.
@@
    We design a multi-teacher on-policy distillation pipeline with mode-aware initialization, which fuses these frontier policies into a single unified student model while preserving strict mode separation and high performance across all modes.
@@ 2 Preliminary
. We study controllable multi-budget reasoning, aiming to learn a single, unified policy $\pi_{\phi}(o\mid q,c)$ that generates distinct, mode-specific reasoning behaviors for a query $q\in\mathcal{D}$, where the mode $c\in\mathcal{C}=\{c_{1},\dots,c_{K}\}$ can be selected or triggered at inference time, corresponding to different reasoning effort.
Crucially, each induced behavior is expected to be strong in its own right: for every mode $c_{k}$, the corresponding policy $\pi_{\phi}(\cdot\mid q,c_{k})$ should achieve high task accuracy relative to its reasoning budget, rather than representing a compromised linear interpolation performance.
@@
Following Ouyang et al. (2022) and Shao et al. (2024), we model language generation as autoregressive sampling from a conditional probability distribution $\pi_{\theta}(y|x)$, and our training goal is to maximize the expected correctness reward $r(q,o)\in\{0,1\}$:
$\max_{\theta}\mathbb{E}_{q,o\sim\pi_{\theta}}[r(q,o)],$ (1)
where $q$ represents the input query and $o$ represents the output response. We adopt RLVR as our training method for two reasons: (i) RLVR provides an effective way to improve reasoning capability by directly reinforcing self-generated sequences with correctness feedback; (ii) In autoregressive models, modifying intermediate tokens in an offline manner—especially in long-form CoT reasoning—can be prone to significant distribution shifts (Xie et al., 2025), whereas on-policy RLVR preserves distributional consistency by training on complete self-generated sequences.
In this work, we adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as our RLVR algorithm for simplicity. Unlike traditional methods (Schulman et al., 2017), GRPO bypasses the value model by computing advantages relative to a group of $G$ responses $\{o_{1},\dots,o_{G}\}$ sampled from $\pi_{\theta_{\text{old}}}$. The objective is:
@@ 3 Methodology
We approach controllable multi-budget reasoning through an exploration–exploitation perspective. The challenge is twofold: for each reasoning mode, the model must achieve high task performance within a specified reasoning budget, while simultaneously supporting multiple modes within a single model.
To address this, we decompose the problem into two complementary phases. In the Exploration phase, multi-stage RL places computational effort in generating sequences in semantic space for exploring under varying context budgets, thus discovering high-quality per-mode strategies. In the Exploitation phase, we consolidate these discovered policies into a single student model via on-policy distillation, preserving mode-specific behaviors while enabling a single efficient model for all reasoning modes. Figure 2 provides an overview of this framework.
@@ 3.1 Expansion-Compression Loop for Mode-Specific Policy Exploration
During exploration, each reasoning mode $c\in\mathcal{C}$ is subject to a context budget reflecting its intended reasoning effort. The training objective is to discover high-quality strategies that maximize task accuracy under this budget. To achieve this, we adopt the simple length regularizing method–truncation along with a scheduling strategy and formalize the exploration stage as traversal along the accuracy–length Pareto frontier.
@@
To efficiently obtain high-quality per-mode strategies under truncation, we introduce a one-time Expansion–Compression Loop for scheduling context window in RLVR training:
@@
    Starting from the expanded policy, we progressively tighten the context budget $L_{k}$. This constraint imposes truncation, inducing a discontinuous length distribution that can destabilize training through two mechanisms: (i) increased reward variance due to the scarcity of valid-length correct trajectories, which leads to noisy gradient estimates; and (ii) “survival mode,” where the model collapses to the shortest, trivial solutions, suppressing the complex reasoning required for harder instances. To balance the need for mode separability with training stability, we adopt a logarithmic compression schedule (i.e., $L_{k+1}=L_{k}/2$). Furthermore, we train each stage to saturation—monitored via a Monte Carlo estimate of the truncation ratio. This ensures that the policy remains locally stable, maintains semantic quality, and preserves a sufficient density of valid trajectories before the next compression.
The loop yields stage-wise, Pareto-optimal per-mode experts along the accuracy–length frontier, which then serve as teachers for on-policy distillation to consolidate all modes into a single model.
@@ 3.2 On-policy Distillation for Frontier Policy Exploitation
While the exploration stage identifies frontier policies, REINFORCE-style updates provide only a sparse $O(1)$ supervision signal per rollout sequence (uniform $\hat{A}$ across all tokens). Such scalar feedback is insufficient not only for efficient consolidation, but also for fine-grained, token-level reasoning behaviors anchoring across multiple mode-specific experts.
[…]
On-Policy Distillation Mode Fusing
Common intuition trains distillation with an offline dataset with query and responses sampled from the teacher model and train the student in an SFT manner.
However, distilling auto-regressive student models in this way leads to a training–inference distribution mismatch, commonly known as exposure bias, because supervision is provided on teacher-generated trajectories that the student policy may never enter during inference, preventing it from learning to correct its own self-induced errors (Agarwal et al., 2023).
Based on these intuition, we adopt On-Policy Distillation (Agarwal et al., 2023; Lu and Lab, 2025), minimizing the Reverse Kullback-Leibler (RKL) divergence:
$\mathcal{L}_{\text{distill}}(\phi)=\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\phi}(\cdot|x,p_{k})}\left[\log\frac{\pi_{\phi}(y|x,p_{k})}{\theta^{(k)}(y|x,p_{k})}\right].$ (11)
RKL is inherently mode-seeking, encouraging the student $\pi_{\phi}$ to focus on the high-density reasoning modes of the expert $\theta^{(k)}$ dictated by prompt $p_{k}$. By concentrating probability mass on these modes, the student is less likely to assign weight to irrelevant trajectories or non-target modes. This ensures that the reasoning knowledge from each expert is effectively consolidated, while preserving the model’s broader generative capabilities.
Formally, the objective function is defined as:
$\mathcal{J}_{\text{Distill}}(\phi)=\mathbb{E}_{k\sim U(1,K)}\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D}\\ o\sim\pi_{\phi}(\cdot|q,p_{k})\end{subarray}}\\ \left[\sum_{t}\log\frac{\pi_{\phi}(o_{t}|q,p_{k},o_{<t})}{\pi_{\theta^{(k)}}(o_{t}|q,o_{<t})}\right],$ (12)
@@
Despite the theoretical advantages of RKL, on-policy distillation suffers a "cold-start": if the student $\pi_{\phi}$ initially lacks confidence in the target modes, rollouts rarely reach the expert’s high-density manifold, limiting the supervisory signal. We mitigate this by initializing the student via model merging, combining the stage-wise frontier models $\{\pi_{\theta^{(k)}}\}_{k=1}^{K}$ obtained during exploration, for aligning input-output mappings with the teacher $\pi_{\theta^{(k)}}$. This provides a more stable start and accelerating convergence under our RKL distillation setting; details can be found in Appendix B.
@@ 4.1 Experimental Settings
Baselines. We first evaluate our full pipeline against the recent controllable reasoning models, including o3-mini (OpenAI, 2025b), gpt-oss-120b (OpenAI, 2025a), and ThinkDial (He et al., 2025). Next, we compare ORBIT’s within-mode performance against current RL-based adaptive and efficient reasoning methods (Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Zhang et al., 2025; Fang et al., 2025; Tu et al., 2025; Liu et al., 2025) on DeepSeek-Distill-Qwen-1.5B. Finally, we ablate ORBIT with other potentially used methods and baselines.
Evaluation Benchmarks. We evaluate our methods on three mathematical reasoning tasks: AIME 2024 & 2025, BeyondAIME (ByteDance-Seed, 2025), as well as representative Multi-task benchmarks: GPQA-Diamond (Rein et al., 2024), and MMLU-Pro (Wang et al., 2024).
@@ 4.2 Main Results
Table 1 illustrates the reasoning behaviors learned by ORBIT under different reasoning budgets across domains. We observe a clear and consistent separation between reasoning modes, with each exhibiting a stable reasoning density and a monotonic trade-off between reasoning effort and task performance. This separation indicates that ORBIT successfully learns query triggered inference-time separable distinct reasoning policies.
@@
ORBIT Achieves Competitive Reasoning Density Within Mode
@@
Table 2: Comparison of reasoning modes of ORBIT trained from DeepSeek-Distill-Qwen-1.5B with efficient reasoning baselines across various reasoning tasks. Darker colors in the cell background denote better results.
We evaluate ORBIT on DeepSeek-Distill-Qwen-1.5B, comparing its within-mode accuracy–length trade-off against RL-based efficient reasoning methods.
As shown in Table 2, ORBIT generally achieves competitive trade-offs, often attaining higher accuracy for a fixed reasoning budget or requiring fewer reasoning tokens for a fixed accuracy level. Notably, these results are obtained with a simple truncation-based reward combined with our scheduling and fusion strategy, highlighting the effectiveness and simplicity of ORBIT.

### v2

@@ Abstract
Recent Large Reasoning Models (LRMs) achieve strong performance by leveraging long-form Chain-of-Thought (CoT) reasoning, but uniformly applying overlong reasoning at inference time incurs substantial and often unnecessary computational cost. To address this, prior work explores various strategies to infer an appropriate reasoning budget from the input. However, such approaches are unreliable in the worst case, as estimating the minimal required reasoning effort is fundamentally difficult, and the trade-off between reasoning cost and accuracy is implicitly fixed during training, limiting flexibility under varying deployment scenarios. Motivated by these limitations, we propose ORBIT, a controllable multi-budget reasoning framework with well-separated reasoning modes triggered by prompt. ORBIT employs multi-stage reinforcement learning to discover Pareto frontier reasoning behaviors at each effort, followed by On-Policy Distillation to fuse these behaviors into a single unified model. Experiments show that ORBIT achieves (1) controllable reasoning behavior over multiple modes, (2) competitive reasoning density within each mode, and (3) integration of these frontier policies into a single unified student model while preserving clear mode separation and token-efficient per-mode performance.
@@ 1 Introduction
Figure 1: ORBIT demonstrates controllable reasoning behaviors comparable to advanced systems, covering multiple reasoning modes: low/mid/high for o3-mini and gpt-oss-120B, and an extended set of low/mid/high/extra-high for ORBIT, illustrating its ability to flexibly trade off reasoning cost and performance across modes.
Recent advances in Reinforcement Learning with Verifiable Rewards (RLVR) (Guo et al., 2025; Team et al., 2025; Xiaomi et al., 2025) have significantly pushed the frontier of complex task solving and long-form Chain-of-Thought (CoT) (Wei et al., 2022) reasoning in large language models. By encouraging explicit reasoning trajectories, these methods enable models to explore richer solution paths and achieve strong performance on challenging benchmarks (Guo et al., 2025). However, the continued push toward greater reasoning depth introduces a practical trade-off between reasoning effort and operational efficiency: increasingly long reasoning traces incur higher latency and computational cost, while often being unnecessary for simpler queries (Yang et al., 2025b).
Notably, this observation concerns not the use of long-form reasoning during training, but its uncalibrated application at inference time, as longer reasoning chains often facilitate exploration of richer solution paths and contribute to higher peak accuracy (Setlur et al., 2025). The challenge primarily emerges at deployment time, where models trained to reason deeply may apply the same heavy reasoning patterns indiscriminately across inputs.
To optimize inference-time efficiency, prior work has explored various strategies for adjusting reasoning effort according to task demands (Luo et al., 2025a; Lou et al., 2025; Luo et al., 2025b; Zhang et al., 2025). Across these approaches, the underlying goal is essentially the same: to estimate an optimal allocation of reasoning effort.
However, this estimation problem is fundamentally difficult. Determining the minimal amount of reasoning required to solve an instance can be viewed as attempting to estimate its Kolmogorov complexity (Li and Vitnyi, 2008), which is known to be uncomputable and connected to the halting problem theoretically (Graves, 2017). Empirically, adaptive mechanisms can also misallocate reasoning effort—truncating reasoning on instances that would benefit from deeper deliberation (Aggarwal et al., 2025; Jin et al., 2025), while preserving unnecessarily long reasoning for simpler instances (Chen et al., 2025a).
@@
Motivated by these limitations, we rethink a controllable reasoning model that can flexibly adjust its reasoning effort according to user preferences by switching reasoning policies across modes at inference time, while achieving competitive reasoning density within each effort level.
We introduce ORBIT, a fully on-policy training framework for controllable multi-budget reasoning. By exploring under different training context windows, ORBIT identifies a Pareto frontier of reasoning behaviors and effectively integrates these into a single model whose discrete reasoning modes can be triggered by specific prompts. As illustrated in Figure 1, ORBIT achieves clear and well-separated reasoning modes similar to state-of-the-art controllable systems, while maintaining strong per-mode accuracy and token-efficient reasoning.
@@
    We propose ORBIT, a fully on-policy training framework for controllable multi-budget reasoning, enabling a single model to generate responses across multiple reasoning modes triggered by mode-specific prompts.
@@
    We introduce an Expansion-Compression Loop exploration strategy that efficiently discovers mode-specific reasoning frontiers under varying context constraints, yielding high quality policies at each reasoning effort.
@@
    We design a Multi-Teacher On-Policy Distillation pipeline with mode-aware initialization, which fuses these frontier policies into a single unified student model while preserving clear mode separation and token-efficient performance across all modes.
@@ 2 Preliminary
We study controllable multi-budget reasoning, aiming to learn a single, unified policy $\pi_{\phi}(o\mid q,c)$ that generates distinct, mode-specific reasoning behaviors for a query $q\in\mathcal{D}$, where the mode $c\in\mathcal{C}=\{c_{1},\dots,c_{K}\}$ can be selected or triggered at inference time, corresponding to different levels of reasoning effort.
Crucially, each induced behavior is expected to be strong in its own right: for every mode $c_{k}$, the conditional policy $\pi_{\phi}(\cdot\mid q,c_{k})$ should achieve high task accuracy relative to its reasoning budget, rather than representing a compromised linear interpolation performance.
@@
We model language generation as autoregressive sampling from a conditional probability distribution $\pi_{\theta}(o|q)$, and our training goal is to maximize the expected correctness reward $r(q,o)\in\{0,1\}$:
$\max_{\theta}\mathbb{E}_{q\sim\mathcal{D},\,o\sim\pi_{\theta}}[r(q,o)],$ (1)
where $q$ represents the input query and $o$ represents the output response.
In this work, we adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as our RLVR algorithm for its simplicity and effectiveness in training reasoning models (Guo et al., 2025). Unlike traditional methods (Schulman et al., 2017), GRPO bypasses the value model by computing advantages relative to a group of $G$ responses $\{o_{1},\dots,o_{G}\}$ sampled from $\pi_{\theta_{\text{old}}}$. The objective is:
@@ 3 Methodology
We approach controllable multi-budget reasoning through an exploration–exploitation perspective. To operationalize, we decompose the problem into two complementary phases:
Exploration Phase. Multi-stage RL allocates computational effort to sample sequences in semantic space under varying context budgets, discovering high-quality per-mode strategies.
Exploitation Phase. Model merging and On-Policy Distillation consolidate these policies into a single student model, preserving mode-specific behaviors while enabling an efficient unified model for all reasoning modes.
Figure 2 provides an overview of this framework.
@@ 3.1 Expansion-Compression Loop for Mode-Specific Policy Exploration
During exploration, each reasoning mode $c\in\mathcal{C}$ is subject to a context budget reflecting its intended reasoning effort. The training objective is to discover high-quality strategies that maximize task accuracy under a given budget. To achieve this, we adopt the simple length regularizing method–truncation, along with a scheduling strategy and formalize the exploration stage as traversal along the accuracy–length Pareto frontier.
@@
To efficiently obtain per-mode strategies under truncation, we introduce a one-time Expansion–Compression Loop for scheduling context window in RLVR training:
@@
    Starting from the expanded policy, we progressively tighten the context budget $L_{k}$. This constraint imposes truncation, inducing a discontinuous length distribution that can destabilize training through two mechanisms: (i) increased reward variance due to the scarcity of valid-length correct trajectories, which leads to noisy gradient estimates; and (ii) “survival mode,” where the model collapses to the shortest, trivial solutions, suppressing the complex reasoning required for harder instances.
    To balance mode separability and training stability, we adopt a logarithmic compression schedule ($L_{k+1}=L_{k}/2$) and sequentially optimize the policy under progressively smaller budgets. This schedule applies comparable truncation pressure on the remaining trajectory mass at each stage, allowing the policy to re-equilibrate under the current constraint before the next compression. Further discussion of this design is provided in Appendix A.
    Notably, binary compression serves as the canonical instantiation of our framework. If finer-grained trade-offs between output quality and latency are desired, additional compression stages can be simply inserted into the current schedule.
The loop yields stage-wise per-mode experts along the accuracy–length frontier, which provide policy anchors in semantic space for consolidating multiple reasoning modes into a single model.
@@ 3.2 On-Policy Distillation for Frontier Policy Exploitation
While the exploration stage identifies frontier policies, REINFORCE-style updates provide only a sparse $O(1)$ supervision signal per rollout sequence (uniform $\hat{A}$ across all tokens). Such scalar feedback is insufficient not only for efficient consolidation, but also for reasoning behaviors anchoring across multiple mode-specific experts.
Transitioning from discovery to consolidation, we therefore adopt distillation as a high-bandwidth alternative, enabling supervision at a tokenwise $O(N)$ resolution. Specifically, we treat the frontier expert models obtained during the exploration stage, $\{\pi_{\theta_{k}}\}_{k=1}^{K}$, as teachers and fuse their mode-specific behaviors into a single unified student model $\pi_{\phi}$. This is achieved through a unified pipeline that combines Multi-Teacher On-Policy Distillation with a Mode-aware Initialization, allowing diverse reasoning modes to be consolidated in a structured and controllable manner.
On-Policy Distillation Mode Fusion
A common intuition is to perform distillation using an offline dataset of query–response pairs sampled from the teacher model, and then train the student in an SFT manner.
However, distilling an autoregressive student model in this way leads to a training–inference distribution mismatch, commonly known as exposure bias (Bengio et al., 2015), because supervision is provided on teacher-generated trajectories that the student policy may never enter during inference, preventing it from learning to calibrate its own behavior on the trajectories induced by its own policy (Agarwal et al., 2023).
Based on these intuitions, we adopt On-Policy Distillation (Agarwal et al., 2023; Lu and Lab, 2025), minimizing the Reverse Kullback-Leibler (RKL) divergence:
$\mathcal{L}_{\text{distill}}(\phi)=\mathbb{E}_{q\sim\mathcal{D},{o}\sim\pi_{\phi}(\cdot|q,p_{k})}\left[\log\frac{\pi_{\phi}({o}|q,p_{k})}{\pi_{\theta_{k}}({o}|q,p_{k})}\right].$ (11)
RKL is inherently mode-seeking, encouraging the student $\pi_{\phi}$ to focus on the high-density reasoning modes of the expert $\pi_{\theta_{k}}$ under prompt $p_{k}$ for mode $c_{k}$. By concentrating probability mass on these modes, the student is less likely to assign weight to irrelevant trajectories or non-target ones. This ensures that the reasoning knowledge from each expert is effectively consolidated, while preserving the model’s broader generative capabilities.
Formally, the objective function is defined as
$\mathcal{L}_{\text{fusion}}(\phi)=\mathbb{E}_{k\sim U(1,K)}\mathbb{E}_{\begin{subarray}{c}q\sim\mathcal{D}\\ {o}\sim\pi_{\phi}(\cdot|q,p_{k})\end{subarray}}\\ \left[\sum_{t}\log\frac{\pi_{\phi}(o_{t}|q,p_{k},o_{<t})}{\pi_{\theta_{k}}(o_{t}|q,p_{k},o_{<t})}\right]$ (12)
@@
Despite the theoretical advantages of RKL, On-Policy Distillation suffers a "cold-start": if the student $\pi_{\phi}$ initially lacks confidence in the target modes, rollouts rarely reach the expert’s high-density manifold, limiting the supervisory signal (Jang et al., 2026). We mitigate this by initializing the student via model merging, combining the stage-wise frontier models $\{\pi_{\theta_{k}}\}_{k=1}^{K}$ obtained during exploration, to align input-output mappings with the teacher $\pi_{\theta_{k}}$. This provides a more stable start and accelerates convergence. Details can be found in Appendix B.
@@ 4.1 Experimental Settings
Baselines. We first evaluate our full pipeline against the recent controllable reasoning models, including o3-mini (OpenAI, 2025), gpt-oss-120b (OpenAI, 2025), and ThinkDial (He et al., 2025). Next, we compare ORBIT’s within-mode performance against current RL-based adaptive and efficient reasoning methods (Aggarwal and Welleck, 2025; Arora and Zanette, 2025; Zhang et al., 2025; Fang et al., 2025; Tu et al., 2025; Liu et al., 2025) on DeepSeek-Distill-Qwen-1.5B. Finally, we ablate ORBIT with other potentially used methods and baselines.
Evaluation Benchmarks. We evaluate our methods on three mathematical reasoning tasks: AIME 2024 & 2025, BeyondAIME (ByteDance-Seed, 2025), as well as representative multi-task benchmarks spanning diverse domains in science, humanities, and professional knowledge: GPQA-Diamond (Rein et al., 2024) and MMLU-Pro (Wang et al., 2024).
@@ 4.2 Main Results
Table 1 illustrates the reasoning behaviors learned by ORBIT under different reasoning budgets across domains. We observe a clear and consistent separation between reasoning modes, with each exhibiting a monotonic trade-off between reasoning effort and task performance. This separation indicates that ORBIT successfully learns prompt triggered inference-time separable distinct reasoning policies.
@@
ORBIT Achieves Competitive Token Efficiency Within Each Mode
@@
Table 2: Comparison of reasoning modes of ORBIT trained from DeepSeek-Distill-Qwen-1.5B with efficient reasoning baselines across various reasoning tasks. Each reasoning mode is evaluated and compared individually.
We evaluate each ORBIT mode separately on DeepSeek-Distill-Qwen-1.5B, comparing its accuracy–length trade-off with RL-based efficient reasoning methods. We aim to determine if the multi-mode student maintains high reasoning density across all triggered efforts without performance compromise.
As shown in Table 2, ORBIT consistently establishes a superior accuracy–length Pareto frontier, reaching higher accuracy under the same reasoning budget or matching baseline accuracy with fewer tokens. Notably, these results are obtained with a minimalist truncation-based reward combined with our scheduling and fusion strategy, highlighting the effectiveness and simplicity of ORBIT.